Skip to end of metadata
Go to start of metadata
Table of Contents


Key:   TPAC1.2 entries in black, TPAC1.1 entries in red, unconfirmed / out of date data in grey.
----------

PCB

Chip

Wfr

Variant

Align

Location

Mods req.

Test summary (current)

History                                           

Comments

 

58

 

 

 

 

 

 

 

Tested 01/12/10 with CoB 1 TP

TPAC Board v2

 

57

 

 

 

 

 

 

 

Tested 01/12/10 with CoB 1 TP

TPAC Board v2

 

56

 

 

 

 

 

 

 

Tested 01/12/10 with CoB 1 & 2 TP

TPAC Board v2

 

55

 

 

 

 

 

 

 

 

TPAC Board v2

 

54

 

 

 

 

 

 

 

11/02/11 SensorLoadv1.1 shows no problems. Testing now
11/02/11 Broken jumper so soldered across pins (test to check performance) TP
Tested 02/12/10 with CoB 1-5 TP

TPAC Board v2

 

50

1.2

14

18u Hires

Center

RAL/JC

0

 

Bonded  22/5/09(W3)
Tested 26/5/09  (config fail)
Bonded 12/06/09 (W3)
Tested 15/06/09
Changed R75=180K 16/6/09
Trimmed at RAL by BL 08/2009
Bonded 15/1/10 (W14)
Trimmed by PD 10/2/10
Bonded 22/2/10 (W14)
Trimmed by PD 22/2/10

 

 

49

1.1

 

12u -DPW 

Laser 

Bristol 

0

 

Bonded  10/2/09 
Tested 19/2/09 

Note this is a TPAC1.1 device!
At Bristol for beam test preparations

 

48

1.2

1

12u +DPW

Center

RAL

0

Trimmed

Bonded 29/5/09
Tested 3/6/09
Trimmed by PD 11/6/09
Removed from 55Fe (dom's lab) 24/7/09
Potted at IC 27/7/09 
Transported to cern (with van) 11/8/09

Rmsd very high in noise scans - may indicate bad trim.
Config errors at 40 degrees in column 37.

 

47

1.2

3

12u +DPW

Center

RAL

0

Trimmed

Bonded 29/5/09(W1)
Tested 3/6/09 (config fail)
Bonded 12/06/09
Tested 15/06/09
Changed R75=180K 16/6/09
Trimmed by PD 23/6/09  
MS/BL report problems with long runs 16/7/09
Potted at RAL 23/7/09
Transported to cern (with van) 11/8/09
Died (no Hits) while at CERN

Died (no Hits) while at CERN

 

46

1.2

3

12u +DPW

Center

RAL/R76

0

Trimmed

Bonded 29/5/09(W1)
Tested 3/6/09 (config fail)
Bonded 12/06/09
Tested 15/06/09
Changed R75=180K 16/6/09
Trimmed by PD 16/6/09
Power module changed 20/7/09
Potted at IC (27/7/09?)
Power shorts found 06/08/09

Running hot, JC suspects some shorting caused by potting.

 

45

1.2

1

12u +DPW

Center

RAL

0

Trimmed
Config errors col 37 only
PD: config errors column 133

Bonded 29/5/09
Tested 3/6/09
Changed R75=180K 11/6/09
Trimmed by PD 15/6/09
Potted at RAL 23/7/09
Transported to cern (with van) 11/8/09
died at CERN (no hits at all)

Died at CERN (no hits at all)   

 

44

1.2

1

12u +DPW

Center

RAL/JC

0

Readback errors region 3
NEW CHIP BONDED

Bonded 26/5/09 (W2)
Tested 29/5/09
Config errors reported by PD 3/6/09
Power module updated to 1.68v 17/7/09
Bonded 26/8/09
Tested 27/08/09

To be reworked with a Hi-Res chip (Dec 09)

 

43

1.2

3

12u +DPW

Center
(TBC)

RAL

0

Trimmed

Bonded 22/5/09
Tested 26/5/09
Tests with 55Fe from 26/05/09
R75=189K 10/7/09
Trimmed by PD ending 17/7/09
Potted at RAL 23/7/09
Transported to cern (with van) 11/8/09

Quick test needed
Note that 43,41,29 are from the first batch of bonding, check which alignment applies!

 

42

1.2

11

12u Hires

Center

RAL/JC

0

Trimmed by PD (feb10) Ok

Bonded 26/5/09(W2)
Tested 29/5/09 (spill fail)
Bonded 12/06/09 (W3)
Tested 15/06/09
Changed R75=180K 16/6/09
Trimmed by PD 30/6/09
DAC register bank errors reported by PD 1/7/09
Adjusted cables, now ok, JC 1/7/09
Changed power module (1.68v) 21/7/09
Bonded 26/8/09 (W1)
Tested 27/08/09
Bonded 15/1/10 (W11)
Trimmed 10/2/10 by PD

Ok

 

 

41

1.2

3

12u +DPW

Center
(TBC)

RAL

0

Trimmed
Config errors col 148 only. PD: see config errors in columns 117 and 129.

Bonded 22/5/09
Tested 26/5/09
Changed R75=180k 1/7/09
Taken to IC by PD 1/7/09
Trimming at IC 2/7/09
Recieved At RAL 27/7/09
Potted at RAL 10/8/09
Carried to cern by JS 12/8/09

Usable
little daughter board came off -> fixed at CERN

 

40

1.1

 

5u +DPW

?

RAL/MS

 

Spill failures reported 28/11/09

Lab9

Ok for analog test pixel measurements

 

39

1.2

3

12u +DPW

Center

RAL

0

To be tested

Bonded 23/7/09
Potted at IC 27/7/09
Trimmed by JC 12/8/09 - problems? TBC
Carried to cern by JS 12/8/09

Analysis of trim runs show last couple are not so good - see trimming page - needs investigation before use!

 

38

1.2

 7

12u -DPW

Center

RAL/JC

 

Good

Bonded 26/8/09
Tested 27/08/09
Trimmed by PD, completed 15/12/09

To be potted

 

37

1.2

11

12u HIRES

Center

RAL

0

To be tested

Bonded 1/7/09
Tested 2/7/09
Power module fitted 2/7/09
R75 changed 2/7/09
Problems in region 0 & 3
Bonded 23/7/09
Potted at IC 27/7/09
Trimming at RAL by JC 07/08/09 
Transported to cern (with van) 11/8/09

Analysis of trim runs show last couple are not so good - see trimming page - needs investigation before use!

 

36

1.2

 

12u -DPW

Center

RAL/JC

 0 

Good

Returned to RAL 1/7/09
Bonded 2/7/09 (W11)
Tested 17/7/09
Trimmed at RAL by BL 08/2009
Bonded 4/11/09 (W7)
Trimmed by PD, completed 11/12/09

To be potted.

 

35

1.2

7

12u -DPW

Center

RAL/JC 

 0

Spill & config load ok but unusually low current (450mA)

Status TBC

Bonded 1/7/09 (Wafer 1 part)
Tested 2/7/09
R75 changed 2/7/09
PD reports no hits 09/07/09
Bonded 23/7/09
Potted at IC 27/7/09
Basic tests at RAL 24/7/09


 

34

1.1

 

5u +DPW 

Laser 

RAL/JC 

 

Ok

 

Keep for laser scans 

 

33

1.2

3

12u +DPW

Center

RAL

 

Low level of data corruption (pixels
appear in data despite being masked)
in region 1, 7/7/09
Trimmed 9/7/09

Bonded 1/7/09
Tested 2/7/09
R75 changed 2/7/09
VDD1V8SRAM testing 6/7/09
Potted at IC 27/7/09
Transported to cern (with van) 11/8/09

 

 

32

1.2

3

12u +DPW

Center

RAL

 

Trimmed. PD: config errors in column 128.

Bonded 1/7/09
Tested 2/7/09
Power module fitted 2/7/09
R75 changed 2/7/09
R84 & R76 changed 2/7/09
Trimmed 3/7/09
PD reported one dead column appeared during trimming 6/7/09
Installed in Dom's Lab (55Fe) 24/7/09
Potted at RAL 10/8/09
Carried to cern by JS 12/8/09

Usable

 

31

1.1

 

5u -DPW

Laser 

RAL/JC 

 

Ok 

 

 

 

30

1.1 

 

12u -DPW 

Laser 

RAL/JC 

 

Ok

 

 

 

29

1.2

3

12u +DPW

Center
(TBC)

RAL

0

Trimmed

Bonded 22/5/09
Tested 26/5/09
Received from IC, 14/7/09
Potted at RAL 10/8/09
Carried to cern by JS 12/8/09

Good

 

28

1.1

 

12u +DPW

Laser 

RAL/JC 

 

Unknown 

 

 Keep for laser scans 

 

27

1.1

 

12u +DPW 

 

RAL/JC 

 

Unknown 

 

Some trailling debug wires
Test & check resistors, mods etc if going to be used

 

26

1.2

14 

18u HIRES

Center

RAL

Trimmed
Pre Beam Test Testing - Errors in columns 3, 9, 20, 25, 27, 38

Returned to RAL 1/7/09
Bonded 2/7/09
Tested 17/7/09
Trimming started 17/7/09 on CALICEDAQ1
Potted at RAL 23/7/09
Transported to cern (with van) 11/8/09

the little power reulgator board come had a broken connection to the middle pin of the regulator (yellow-green) , re-soldered

 

25

 

 

 

 

 

 

 

 

 

 

24

1.1

 

12u +DPW

 

RAL/MS

 

 

Lab9

 

 

23

1.2

13

18u HIRES

Laser

RAL/MS

0

Ok

Bonded 29/5/09
Tested 3/6/09
Changed R75=180K 11/6/09
test with 55Fe from 18/6/09

 

 

22

-

-

-

-

RAL/JC

-

Bare-PCB Test shows faulty DAC bank

Bare board test 6/5/09 failed

 

 

21

1.2

9

12u HIRES

Laser

RAL

0

Config errors col 148 only

Bonded 26/5/09
Tested 29/5/09
Tests with 55Fe from 1/6/09
Changed R75=180K 27/7/09
Trimming started 27/7/09
Potted at RAL 10/8/09
Carried to cern by JS 12/8/09

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 



 

* Mods indicates the number of outstanding pcb modifications required to bring this PCB up to date.  A full list of PCB mods is shown here

  • No labels